# Generalization Error of the Tilted Empirical Risk

### v1

@@ 1 Introduction
A central concern in statistical learning theory is understanding the efficacy of a learning algorithm when applied to test data. This evaluation is typically carried out by investigating the generalization error, which quantifies the disparity between the performance of the algorithm on the training dataset and its performance on previously unseen data, drawn from the same underlying distribution, via a risk function. Understanding the generalization behaviour of learning algorithms is one of the most important objectives in statistical learning theory. Various approaches have been developed (Rodrigues and Eldar, 2021), including VC dimension-based bounds (Vapnik, 1999), stability-based bounds (Bousquet and Elisseeff, 2002b), PAC Bayesian bounds (McAllester, 2003), and information-theoretic bounds (Russo and Zou, 2019; Xu and Raginsky, 2017). However, the TER generalization performance has not yet been studied. This paper focuses on the generalization error of the tilted empirical risk (tilted generalization error) of learning algorithms. Our contributions here can be summarized as follows,
  • We provide upper bounds on the tilted generalization error via uniform and information-theoretic approaches under bounded loss functions for both positive and negative tilts and show that the convergence rates of the upper bounds are $O(1/\sqrt{n})$, as expected, where $n$ is the number of training samples.
  • We provide upper bounds on the tilted generalization error under unbounded loss functions for the negative tilt with the convergence rate of $O(1/\sqrt{n})$ via a uniform approach and and information-theoretic approach. Our tools could be used to establish convergence of the empirical risk for unbounded losses.
@@
The paper is organised as follows: Section 2 introduces notation, the problem, and the risk functions used in this paper. Our upper and lower bounds on the tilted generalization error for bounded loss functions via uniform and information-theoretic approaches are given in Section 3. The upper bounds on unbounded loss functions with bounded second moments are given in Section 4. Section 5 is devoted to the study of robustness to the distributional shift in training samples. The KL-regularized TERM is considered in Section 6. Section 7 surveys related work. Section 8 concludes the paper. Technical tools and proofs are deferred to the appendices.
@@ 2.2 Risk Functions
$\begin{split}\mathrm{R}(H,P_{H}\otimes\mu):=\mathbb{E}_{P_{H}\otimes\mu}[\ell(H,Z)],&\qquad\overline{\mathrm{R}}_{\gamma}(H,P_{H,S}):=\mathbb{E}_{P_{H,S}}[\widehat{\mathrm{R}}_{\gamma}(H,S)],\\ \overline{\mathrm{gen}}_{\gamma}(H,S)&:=\mathbb{E}_{P_{H,S}}[\mathrm{gen}_{\gamma}(H,S)].\end{split}$ (11)
Similar to (11), we define the tilted population risk and the tilted generalization error, (9) and (10), as expectations. We also provide upper bounds on the expected tilted generalization error with respect to the joint distribution of $S$ and $H$, of the form
$\overline{\mathrm{gen}}_{\gamma}(H,S)\leq g_{e}(n,\gamma),$
where $g_{e}$ is a real function. An overview of our main results is provided in Fig. 1.
Tilted Generalization Error Bounded Loss Function ($\gamma\in\mathbb{R}$) Section 3 Information-theoretic Bound Upper bound (Theorem 3.13) & Lower bound (Theorem 3.14) Uniform Bound Upper bound (Theorem 3.3) & Lower bound (Theorem 3.4) Unbounded Loss Function ($\gamma\in\mathbb{R}_{<0}$) Section 4 Information-theoretic Bound Upper bound (Theorem 4.12) & Lower bound (Theorem 4.14) Uniform Bound Upper bound (Theorem 4.5) & Lower bound (Theorem 4.6) Robustness of TERM ($\gamma\in\mathbb{R}_{<0}$) Section 5 Uniform Bound Upper bound (Theorem 5.4) KL-Regularized TERM ($\gamma\in\mathbb{R}$) Section 6 Information-theoretic Bound Upper bound (Theorem 6.3)
Figure 1: Overview of the main results
@@ 3 Generalization Bounds for Bounded Loss Function
Upper bounds under linear empirical risk for bounded loss functions via information theoretic and uniform bounds are studied by Shalev-Shwartz and Ben-David (2014) and Xu and Mannor (2012), respectively. Inspired by these works, in this section, we provide upper bounds on the tilted generalization error via uniform and information-theoretic approaches for bounded loss functions with the convergence rate of $O(1/\sqrt{n})$ which is similar to generalization error under linear empirical risk. Upper bounds via stability (Bousquet and Elisseeff, 2002b), Rademacher complexity (Bartlett and Mendelson, 2002) and PAC-Bayesian approaches (Alquier, 2021) are provided in Appendix G. All proof details are deferred to Appendix C.
@@ 3.1 Uniform Bounds
$\begin{split}&\mathrm{gen}_{\gamma}(h,S)=\underbrace{\mathrm{R}(h,\mu)-\mathrm{R}_{\gamma}(h,\mu^{\otimes n})}_{I_{1}}+\underbrace{\mathrm{R}_{\gamma}(h,\mu^{\otimes n})-\widehat{\mathrm{R}}_{\gamma}(h,S)}_{I_{2}},\end{split}$ (12)
@@
$\begin{split}\frac{-1}{2\gamma}\mathrm{Var}\big{(}\exp(\gamma\ell(h,Z))\big{)}\leq\mathrm{R}(h,\mu)-\mathrm{R}_{\gamma}(h,\mu^{\otimes n})\leq\frac{-\exp(-2\gamma M)}{2\gamma}\mathrm{Var}\big{(}\exp(\gamma\ell(h,Z))\big{)}.\end{split}$ (13)
@@
$\begin{split}\mathrm{gen}_{\gamma}(h,S)\leq\frac{-\exp(-2\gamma M)}{2\gamma}\operatorname{Var}\big{(}\exp(\gamma\ell(h,Z))\big{)}+\frac{\big{(}\exp(|\gamma|M)-1\big{)}}{|\gamma|}\sqrt{\frac{\log(2/\delta)}{2n}}.\end{split}$ (14)
@@
$\begin{split}\mathrm{gen}_{\gamma}(h,S)\geq\frac{-1}{2\gamma}\operatorname{Var}\big{(}\exp(\gamma\ell(h,Z))\big{)}-\frac{\big{(}\exp(|\gamma|M)-1\big{)}}{|\gamma|}\sqrt{\frac{\log(2/\delta)}{2n}}.\end{split}$ (15)
@@
The lower bound in Theorem 3.4 for a negative $\gamma$ can be tighter than the lower bound on the tilted generalization error, (15), as the variance is positive.
Remark 3.6.
@@
$\begin{split}\sup_{h\in\mathcal{H}}|\mathrm{gen}_{\gamma}(h,S)|&\leq\frac{\big{(}\exp(|\gamma|M)-1\big{)}}{|\gamma|}\sqrt{\frac{\log(\mathrm{card}(\mathcal{H}))+\log(2/\delta)}{2n}}+\frac{\max(1,\exp(-2\gamma M))A(\gamma)}{8|\gamma|},\end{split}$
@@
$\begin{split}\mathfrak{E}_{\gamma}(\mu)&\leq\frac{2\big{(}\exp(|\gamma|M)-1\big{)}}{|\gamma|}\sqrt{\frac{\log(\mathrm{card}(\mathcal{H}))+\log(2/\delta)}{2n}}+\frac{2\max(1,\exp(-2\gamma M))A(\gamma)}{8|\gamma|},\end{split}$
@@ 3.2 Information-theoretic Bounds
Next, we provide an upper bound on the expected tilted generalization error. All proof details are deferred to Appendix C.2. For information-theoretic bounds, we employ the following decomposition of the expected tilted generalization error,
$\begin{split}&\overline{\mathrm{gen}}_{\gamma}(H,S)=\{\mathrm{R}(H,P_{H}\otimes\mu)-\overline{\mathrm{R}}_{\gamma}(H,P_{H}\otimes\mu^{\otimes n})\}+\{\overline{\mathrm{R}}_{\gamma}(H,P_{H}\otimes\mu^{\otimes n})-\overline{\mathrm{R}}_{\gamma}(H,P_{H,S})\}.\end{split}$ (17)
@@
$\begin{split}&\Big{|}\overline{\mathrm{R}}_{\gamma}(H,P_{H}\otimes\mu^{\otimes n})-\overline{\mathrm{R}}_{\gamma}(H,P_{H,S})\Big{|}\leq\frac{(\exp(|\gamma|M)-1)}{|\gamma|}\sqrt{\frac{I(H;S)}{2n}}.\end{split}$ (18)
@@
$\overline{\mathrm{gen}}_{\gamma}(H,S)\geq-\frac{(\exp(|\gamma|M)-1)}{|\gamma|}\sqrt{\frac{I(H;S)}{2n}}-\frac{\gamma\exp(\gamma M)}{2}\Big{(}1-\frac{1}{n}\Big{)}\mathbb{E}_{P_{H}}\big{[}\mathrm{Var}_{\tilde{Z}\sim\mu}(\ell(H,\tilde{Z}))\big{]}.$
@@
$\begin{split}|\overline{\mathrm{gen}}_{\gamma}(H,S)|&\leq\frac{(\exp(|\gamma|M)-1)}{|\gamma|}\sqrt{\frac{I(H;S)}{2n}}+\frac{|\gamma|M^{2}\exp(|\gamma|M)}{8}\Big{(}1-\frac{1}{n}\Big{)}.\end{split}$
@@
In Theorem 3.13, we observe that by choosing $\gamma=O(n^{-\beta})$, the overall convergence rate of the generalization error upper bound is $\max(O(1/\sqrt{n}),O(n^{-\beta}))$ for bounded $I(H;S)$. For $\beta\geq 1/2$, the convergence rate of (19) is the same as the convergence rate of the expected upper bound in (Xu and Raginsky, 2017) . In addition, for $\gamma\rightarrow 0$, the upper bound in Theorem 3.13 converges to the expected upper bound in (Xu and Raginsky, 2017).
@@ 4 Generalization Bounds for Unbounded Loss Functions
In the previous section, we assumed that the loss function is bounded which is limiting in practice. Several works already proposed some solutions to overcome the boundedness assumption under linear empirical risk (Lugosi and Neu, 2022; Haddouche and Guedj, 2022; Alquier and Guedj, 2018; Holland, 2019) via a PAC-Bayesian approach. In this section, we derive upper bounds on the tilted generalization error via uniform and information-theoretic approaches for negative tilt ($\gamma<0$) under bounded second moment assumption with convergence rate of $O(1/\sqrt{n})$. In particular, we relax the bounded assumption (Assumption 3.1). The following assumptions are made for the uniform analysis.
Assumption 4.1 (Uniform bounded true risk).
There is a constant $R_{T}^{u}$ such that the loss function $\ell$ satisfies $\mathbb{E}_{\mu}[\ell(h,\tilde{Z})]\leq R_{T}^{u}<\infty$ uniformly for all $h\in\mathcal{H}$.
@@
There is a constant $M_{2}^{u}$ such that the loss function $(H,Z)\mapsto\ell(H,Z)$ satisfies $\mathbb{E}_{\mu}[\ell^{2}(h,Z)]\leq M_{2}^{u}$ uniformly for all $h\in\mathcal{H}$.
In the information-theoretic approach for the unbounded loss function, we relax the uniform assumption, Assumptions 4.1 and 4.2, as follows,
Assumption 4.3 (Bounded true risk).
The learning algorithm $P_{H|S}$, loss function $\ell$, and $\mu$ are such that the expected population risk satisfies $\mathbb{E}_{P_{H}\otimes\mu}[\ell(H,\tilde{Z})]=R_{T}<\infty$.
Assumption 4.4 (Bounded second moment).
The learning algorithm $P_{H|S}$, loss function $\ell$, and $\mu$ are such that there is a constant $M_{2}$ where the loss function $(H,Z)\mapsto\ell(H,Z)$ satisfies $\max\big{(}\mathbb{E}_{P_{H,Z}}[\ell^{2}(H,Z)],\mathbb{E}_{P_{H}\otimes\mu}[\ell^{2}(H,Z)]\big{)}\leq M_{2}$ for all $Z\in\mathcal{Z}$.
The assumptions on second moments, Assumption 4.2 and 4.4, can be satisfied if the loss function is sub-Gaussian or sub-Exponential (Boucheron et al., 2013) under the distribution $\mu$ for all $h\in\mathcal{H}$. All proof details for the results in this section are deferred to Appendix D.
4.1 Uniform Bounds
For unbounded loss function, we consider the decomposition of the tilted generalization error in (12) for $\gamma<0$. The term $I_{1}$ can be bounded using Lemma B.7. Then, we apply Bernstein’s inequality (Boucheron et al., 2013) to provide upper and lower bounds on the second term, $I_{2}$.
@@ 4.1 Uniform Bounds
Given Assumption 4.1 and Assumption 4.2, for any fixed $h\in\mathcal{H}$ with probability at least $(1-\delta)$, then the following upper bound holds on the tilted generalization error for $\gamma<0$,
$\begin{split}\mathrm{gen}_{\gamma}(h,S)&\leq 2\exp(-\gamma R_{T}^{u})\sqrt{\frac{M_{2}^{u}\log(2/\delta)}{n}}-\frac{4\exp(-\gamma R_{T}^{u})\log(2/\delta)}{3n\gamma}-\frac{\gamma}{2}M_{2}^{u}.\end{split}$ (20)
@@
Given Assumption 4.1 and Assumption 4.2, there exists a $\zeta\in(0,1)$ such that for $n\geq\frac{(4\gamma^{2}M_{2}^{u}+8/3\zeta)\log(2/\delta)}{\zeta^{2}\exp(2\gamma R_{T}^{u})}$, for any fixed $h\in\mathcal{H}$ with probability at least $(1-\delta)$, and $\gamma<0$, the following lower bound on the tilted generalization error holds
$\mathrm{gen}_{\gamma}(h,S)\geq-\frac{2\exp(-\gamma R_{T}^{u})}{(1-\zeta)}\sqrt{\frac{M_{2}^{u}\log(2/\delta)}{n}}+\frac{4\exp(-\gamma R_{T}^{u})(\log(2/\delta))}{3n\gamma(1-\zeta)}.$ (21)
@@
Under the same assumptions in Theorem 4.6 and a finite hypothesis space, then for $n\geq\frac{(4\gamma^{2}M_{2}^{u}+8/3\zeta)\log(2/\delta)}{\zeta^{2}\exp(2\gamma R_{T}^{u})}$, for $\gamma<0$ and with probability at least $(1-\delta)$, the absolute value of the titled generalization error satisfies
$\begin{split}\sup_{h\in\mathcal{H}}|\mathrm{gen}_{\gamma}(h,S)|&\leq\frac{2\exp(-\gamma R_{T}^{u})}{(1-\zeta)}\sqrt{\frac{M_{2}^{u}(\log(\mathrm{card}(\mathcal{H}))+\log(2/\delta))}{n}}\\ &\qquad-\frac{4\exp(-\gamma R_{T}^{u})(\log(\mathrm{card}(\mathcal{H}))+\log(2/\delta))}{3n\gamma(1-\zeta)}-\frac{\gamma}{2}M_{2}^{u}.\end{split}$ (22)
@@
For $\gamma\asymp n^{-1/2}$, the upper bound in Corollary 4.7 gives a theoretical guarantee on the convergence rate of $O(n^{-1/2})$. Using TER with negative $\gamma$ can help to derive an upper bound on the absolute value of the tilted generalization error under the bounded second-moment assumption.
Similar to Corollary 3.11, an upper bound on excess risk under unbounded loss function assumptions can be derived.

### v2

@@ 1 Introduction
A central concern in statistical learning theory is understanding the efficacy of a learning algorithm when applied to test data. This evaluation is typically carried out by investigating the generalization error, which quantifies the disparity between the performance of the algorithm on the training dataset and its performance on previously unseen data, drawn from the same underlying distribution, via a risk function. Understanding the generalization behaviour of learning algorithms is one of the most important objectives in statistical learning theory. Various approaches have been developed (Rodrigues and Eldar, 2021), including VC dimension-based bounds (Vapnik, 1999), stability-based bounds (Bousquet and Elisseeff, 2002b), PAC Bayesian bounds (McAllester, 2003), and information-theoretic bounds (Russo and Zou, 2019; Xu and Raginsky, 2017). However, the TER generalization performance under unbounded loss function has not yet been studied. This paper focuses on the generalization error of the tilted empirical risk (tilted generalization error) of learning algorithms. Our contributions here can be summarized as follows,
  • We provide upper and lower bounds on the tilted generalization error via uniform and information-theoretic approaches under bounded loss functions for both positive and negative tilts and show that the convergence rates of the upper bounds are $O(1/\sqrt{n})$, as expected, where $n$ is the number of training samples.
  • We provide upper and lower bounds on the tilted generalization error under unbounded loss functions for the negative tilt with the convergence rate of $O(1/\sqrt{n})$ via a uniform approach.
@@
The paper is organised as follows: Section 2 introduces notation, the problem, and the risk functions used in this paper. Our upper and lower bounds on the tilted generalization error for bounded loss functions via uniform and information-theoretic approaches are given in Section 3. The upper and lower bounds on the tilted generalization error for unbounded loss functions with bounded second moments are given in Section 4. Section 5 is devoted to the study of robustness to the distributional shift in training samples. The KL-regularized TERM is considered in Section 6. Section 7 surveys related work. Section 8 concludes the paper. Technical tools and proofs are deferred to the appendices.
@@ 2.2 Risk Functions
$\begin{split}\mathrm{R}(H,P_{H}\otimes\mu)&:=\mathbb{E}_{P_{H}\otimes\mu}[\ell(H,Z)],\\ \overline{\mathrm{R}}_{\gamma}(H,P_{H,S})&:=\mathbb{E}_{P_{H,S}}[\widehat{\mathrm{R}}_{\gamma}(H,S)],\\ \overline{\mathrm{gen}}_{\gamma}(H,S)&:=\mathbb{E}_{P_{H,S}}[\mathrm{gen}_{\gamma}(H,S)].\end{split}$ (11)
Similar to (11), we define the tilted population risk and the tilted generalization error, (9) and (10), as expectations. We also provide upper bounds on the absolute value of expected tilted generalization error with respect to the joint distribution of $S$ and $H$, of the form
$|\overline{\mathrm{gen}}_{\gamma}(H,S)|\leq g_{e}(n,\gamma),$
where $g_{e}$ is a real function.
@@ 3 Generalization Bounds for Bounded Loss Function
Upper bounds under linear empirical risk for bounded loss functions via information theoretic and uniform bounds are studied by Shalev-Shwartz and Ben-David (2014) and Xu and Mannor (2012), respectively. Inspired by these works, in this section, we provide upper bounds on the tilted generalization error via uniform and information-theoretic approaches for bounded loss functions with the convergence rate of $O(1/\sqrt{n})$ which is similar to generalization error under linear empirical risk. Upper bounds via stability (Bousquet and Elisseeff, 2002b), Rademacher complexity (Bartlett and Mendelson, 2002) and PAC-Bayesian approaches (Alquier, 2021) are provided in Appendix H. All proof details are deferred to Appendix D.
@@ 3.1 Uniform Bounds
$\begin{split}&\mathrm{gen}_{\gamma}(h,S)=\underbrace{\mathrm{R}(h,\mu)-\mathrm{R}_{\gamma}(h,\mu^{\otimes n})}_{I_{1}}\\ &\qquad+\underbrace{\mathrm{R}_{\gamma}(h,\mu^{\otimes n})-\widehat{\mathrm{R}}_{\gamma}(h,S)}_{I_{2}},\end{split}$ (12)
@@
$\begin{split}&\frac{-1}{2\gamma}\mathrm{Var}\big{(}\exp(\gamma\ell(h,Z))\big{)}\leq\mathrm{R}(h,\mu)-\mathrm{R}_{\gamma}(h,\mu^{\otimes n})\\ &\leq\frac{-\exp(-2\gamma M)}{2\gamma}\mathrm{Var}\big{(}\exp(\gamma\ell(h,Z))\big{)}.\end{split}$ (13)
@@
$\begin{split}\mathrm{gen}_{\gamma}(h,S)&\leq\frac{-\exp(-2\gamma M)}{2\gamma}\operatorname{Var}\big{(}\exp(\gamma\ell(h,Z))\big{)}\\ &\qquad+\frac{\big{(}\exp(|\gamma|M)-1\big{)}}{|\gamma|}\sqrt{\frac{\log(2/\delta)}{2n}}.\end{split}$
@@
$\begin{split}\mathrm{gen}_{\gamma}(h,S)&\geq\frac{-1}{2\gamma}\operatorname{Var}\big{(}\exp(\gamma\ell(h,Z))\big{)}\\ &\qquad-\frac{\big{(}\exp(|\gamma|M)-1\big{)}}{|\gamma|}\sqrt{\frac{\log(2/\delta)}{2n}}.\end{split}$ (14)
@@
The lower bound in Theorem 3.4 for a negative $\gamma$ can be tighter than the lower bound on the tilted generalization error for a positive $\gamma$, (14), as the variance is positive.
@@
$\begin{split}&\sup_{h\in\mathcal{H}}|\mathrm{gen}_{\gamma}(h,S)|\\ &\leq\frac{\big{(}\exp(|\gamma|M)-1\big{)}}{|\gamma|}\sqrt{\frac{\log(\mathrm{card}(\mathcal{H}))+\log(2/\delta)}{2n}}\\ &\qquad+\frac{\max(1,\exp(-2\gamma M))A(\gamma)}{8|\gamma|},\end{split}$
@@
$\begin{split}\mathfrak{E}_{\gamma}(\mu)&\leq\frac{2\big{(}\exp(|\gamma|M)-1\big{)}}{|\gamma|}\sqrt{\frac{\log(\mathrm{card}(\mathcal{H}))+\log(2/\delta)}{2n}}\\ &\qquad+\frac{2\max(1,\exp(-2\gamma M))A(\gamma)}{8|\gamma|},\end{split}$
@@ 3.2 Information-theoretic Bounds
Next, we provide an upper bound on the expected tilted generalization error. All proof details are deferred to Appendix D.2. For information-theoretic bounds, we employ the following decomposition of the expected tilted generalization error,
$\begin{split}&\overline{\mathrm{gen}}_{\gamma}(H,S)\\ &=\{\mathrm{R}(H,P_{H}\otimes\mu)-\overline{\mathrm{R}}_{\gamma}(H,P_{H}\otimes\mu^{\otimes n})\}\\ &+\{\overline{\mathrm{R}}_{\gamma}(H,P_{H}\otimes\mu^{\otimes n})-\overline{\mathrm{R}}_{\gamma}(H,P_{H,S})\}.\end{split}$ (16)
@@
$\begin{split}&\Big{|}\overline{\mathrm{R}}_{\gamma}(H,P_{H}\otimes\mu^{\otimes n})-\overline{\mathrm{R}}_{\gamma}(H,P_{H,S})\Big{|}\\ &\leq\frac{(\exp(|\gamma|M)-1)}{|\gamma|}\sqrt{\frac{I(H;S)}{2n}}.\end{split}$ (17)
@@
$\begin{split}&\overline{\mathrm{gen}}_{\gamma}(H,S)\geq-\frac{(\exp(|\gamma|M)-1)}{|\gamma|}\sqrt{\frac{I(H;S)}{2n}}\\ &\qquad-\frac{\gamma\exp(\gamma M)}{2}\Big{(}1-\frac{1}{n}\Big{)}\mathbb{E}_{P_{H}}\big{[}\mathrm{Var}_{\tilde{Z}\sim\mu}(\ell(H,\tilde{Z}))\big{]}.\end{split}$
@@
$\begin{split}|\overline{\mathrm{gen}}_{\gamma}(H,S)|&\leq\frac{(\exp(|\gamma|M)-1)}{|\gamma|}\sqrt{\frac{I(H;S)}{2n}}\\ &\qquad+\frac{|\gamma|M^{2}\exp(|\gamma|M)}{8}\Big{(}1-\frac{1}{n}\Big{)}.\end{split}$
@@
In Corollary 3.15, we observe that by choosing $\gamma=O(n^{-\beta})$, the overall convergence rate of the generalization error upper bound is $\max(O(1/\sqrt{n}),O(n^{-\beta}))$ for bounded $I(H;S)$. For $\beta\geq 1/2$, the convergence rate of (18) is the same as the convergence rate of the expected upper bound in (Xu and Raginsky, 2017). In addition, for $\gamma\rightarrow 0$, the upper bound in Corollary 3.15 converges to the expected upper bound in (Xu and Raginsky, 2017).
The results in this section are non-vacuous for bounded $I(H;S)$. If this assumption is violated, we can apply the individual sample method (Bu et al., 2020), chaining methods (Asadi et al., 2018), or conditional mutual information frameworks (Steinke and Zakynthinou, 2020) to derive tighter upper bound for the tilted generalization error.
@@ 4 Generalization Bounds for Unbounded Loss Functions
In the previous section, we assumed that the loss function is bounded which is limiting in practice. Several works already proposed some solutions to overcome the boundedness assumption under linear empirical risk (Lugosi and Neu, 2022; Haddouche and Guedj, 2022; Alquier and Guedj, 2018; Holland, 2019) via a PAC-Bayesian approach. In this section, we derive upper bounds on the tilted generalization error via uniform approach for negative tilt ($\gamma<0$) under bounded second moment assumption with convergence rate of $O(1/\sqrt{n})$. In particular, we relax the bounded assumption (Assumption 3.1). The following assumptions are made for the uniform analysis.
@@
There is a constant $\kappa_{u}\in\mathbb{R}^{+}$ such that the loss function $(H,Z)\mapsto\ell(H,Z)$ satisfies $\mathbb{E}_{\mu}[\ell^{2}(h,Z)]\leq\kappa_{u}^{2}$ uniformly for all $h\in\mathcal{H}$.
The assumption on second moment, Assumption 4.1, can be satisfied if the loss function is sub-Gaussian or sub-Exponential (Boucheron et al., 2013) under the distribution $\mu$ for all $h\in\mathcal{H}$. All proof details for the results in this section are deferred to Appendix E. In this following we present the results based on uniform bounds. The results for unbounded loss function based on information theoretic bounds are presented in Appendix E.2.
For unbounded loss function, we consider the decomposition of the tilted generalization error in (12) for $\gamma<0$. The term $I_{1}$ can be bounded using Lemma C.9. Then, we apply Bernstein’s inequality (Boucheron et al., 2013) to provide upper and lower bounds on the second term, $I_{2}$.
@@
Given Assumption 4.1, for any fixed $h\in\mathcal{H}$ with probability at least $(1-\delta)$, then the following upper bound holds on the tilted generalization error for $\gamma<0$,
$\begin{split}\mathrm{gen}_{\gamma}(h,S)&\leq 2\kappa_{u}\exp(-\gamma\kappa_{u})\sqrt{\frac{\log(2/\delta)}{n}}\\ &-\frac{4\exp(-\gamma\kappa_{u})\log(2/\delta)}{3n\gamma}-\frac{\gamma}{2}\kappa_{u}^{2}.\end{split}$ (19)
@@
Given Assumption 4.1, there exists a $\zeta\in(0,1)$ such that for $n\geq\frac{(4\gamma^{2}\kappa_{u}^{2}+8/3\zeta)\log(2/\delta)}{\zeta^{2}\exp(2\gamma\kappa_{u})}$, for any fixed $h\in\mathcal{H}$ with probability at least $(1-\delta)$, and $\gamma<0$, the following lower bound on the tilted generalization error holds
$\begin{split}\mathrm{gen}_{\gamma}(h,S)&\geq-\frac{2\kappa_{u}\exp(-\gamma\kappa_{u})}{(1-\zeta)}\sqrt{\frac{\log(2/\delta)}{n}}\\ &+\frac{4\exp(-\gamma\kappa_{u})(\log(2/\delta))}{3n\gamma(1-\zeta)}.\end{split}$ (20)
@@
Under the same assumptions in Theorem 4.3 and a finite hypothesis space, then for $n\geq\frac{(4\gamma^{2}\kappa_{u}^{2}+8/3\zeta)\log(2/\delta)}{\zeta^{2}\exp(2\gamma\kappa_{u})}$, for $\gamma<0$ and with probability at least $(1-\delta)$, the absolute value of the titled generalization error satisfies
$\sup_{h\in\mathcal{H}}|\mathrm{gen}_{\gamma}(h,S)|\leq\frac{2\kappa_{u}\exp(-\gamma\kappa_{u})}{(1-\zeta)}\sqrt{\frac{B(\delta)}{n}}-\frac{4\exp(-\gamma\kappa_{u})B(\delta)}{3n\gamma(1-\zeta)}-\frac{\gamma}{2}\kappa_{u}^{2},$ (21)
where $B(\delta)=\log(\mathrm{card}(\mathcal{H}))+\log(2/\delta)$.
@@
For $\gamma\asymp n^{-1/2}$, $n\geq\frac{(4\gamma^{2}\kappa_{u}^{2}+8/3\zeta)\log(2/\delta)}{\zeta^{2}\exp(-2\kappa_{u})}$ and $-1<\gamma<0$, the upper bound in Corollary 4.4 gives a theoretical guarantee on the convergence rate of $O(n^{-1/2})$. Using TER with negative $\gamma$ can help to derive an upper bound on the absolute value of the tilted generalization error under the bounded second-moment assumption.
Similar to Corollary 3.11, an upper bound on excess risk under unbounded loss function assumption can be derived.